# A Smooshed BMOCZ Zero Constellation for CFO Estimation Without Channel Coding

### v1

@@ III-B CFO Correction with SBMOCZ
In this subsection, we introduce a method to estimate the CFO by evaluating the rotated polynomial $\tilde{Y}(z)$ at various points on the unit circle. To justify evaluating on the unit circle, we highlight some useful properties. We begin with the following proposition:
@@
Let $X(z)$ be a polynomial with coefficients $\boldsymbol{\rm{x}}=(x_{0},x_{1},\dots,x_{K})\in\mathbb{C}^{K+1}$. Then, the squared magnitude of $X(z)$ evaluated at $z=e^{-j\theta}$ for $\theta\in\left[0,2\pi\right)$ corresponds to the discrete-time Fourier transform (DTFT) of its auto-correlation sequence $\mathbb{a}=(a_{-K},a_{-K+1},\dots,a_{K})\in\mathbb{C}^{2K+1}$. Specifically, we have
$\left|X(e^{-j\theta})\right|^{2}=\sum_{\ell=-K}^{K}a_{\ell}e^{-j\theta\ell}\;.$ (10)
@@
The auto-correlation sequence $\mathbb{a}$ is the same for any transmitted BMOCZ sequence $\boldsymbol{\rm{x}}$ [4]. Consequently, the evaluation $|X(e^{-j\theta})|^{2}$ yields the same result for any SBMOCZ polynomial.
The important takeaway from Proposition 1 and Corollary 1 is that the zero constellation gap will induce a peak on the unit circle at the same location for any transmitted SBMOCZ sequence $\boldsymbol{\rm{x}}$. Additionally, evaluation on the unit circle will be symmetric about the peak, as highlighted by the following lemma:
@@
Let $X(z)$ be any SBMOCZ polynomial. When evaluating $|X(e^{-j\theta})|^{2}$ for $\theta\in[0,2\pi)$, we have
$\left|X(e^{-j\theta})\right|^{2}=\left|X(e^{j\theta})\right|^{2}\;.$ (11)
The proof is given in Appendix B. For SBMOCZ, the gap is centered on the positive real axis, resulting in a peak on the unit circle at $z=1$. To illustrate this behavior, we now present Fig. 2, which plots $|X(z)|^{2}$ along the unit circle for different $\zeta$ with $K=32$. As the gap increases in size, the peak becomes more pronounced, which makes the CFO estimate more robust against noise. However, this comes at the cost of larger zero-perturbation under noise. Additionally, Fig. 2 shows that Huffman BMOCZ is incompatible with our approach, since there is not a unique maximum on $\left[0,2\pi\right)$. Instead, evaluation on the unit circle exhibits sinusoidal behavior for Huffman BMOCZ. The key observation is that by identifying the peak location for $|\tilde{Y}(z)|^{2}$ on the unit circle, we can obtain an estimate for the angular rotation $\psi$.
Without loss of generality, we will now consider $|\tilde{Y}(z)|$, which peaks at the same location as its square. Under rotation, the maximum of $|X(e^{j(\theta+\psi)})|$ in SBMOCZ occurs when $\theta=-\psi$, so we can estimate the angular rotation from $\tilde{Y}(z)$ via
$\hat{\psi}=\arg\!\!\!\,\max_{\theta\in\left[0,2\pi\right)}\left|\tilde{Y}(e^{-j\theta})\right|\;.$ (12)
@@
$\hat{\psi}\approx\frac{2\pi}{N}\arg\max_{n\in[N]}\left|\tilde{Y}(e^{-j2\pi n/N})\right|\>,$ (13)
where $N$ determines the resolution of the search space. We highlight that $\tilde{Y}(e^{-j2\pi n/N})$ can be evaluated using a single $N$-point DFT of the coefficients, which results in a modest time complexity of $\mathcal{O}(N\log N)$ for the proposed algorithm. To compensate for the rotation, we perform the correction $\hat{Y}(z)=\tilde{Y}(e^{-j\hat{\psi}}z)$, where $\hat{Y}(z)$ is an estimate of the received polynomial without CFO impairment. The DiZeT decoder is then applied to $\hat{Y}(z)$ to recover the binary message $\boldsymbol{\rm{m}}$.
@@ IV Numerical results
In this section, we compare the performance of SBMOCZ to Huffman BMOCZ in AWGN and flat-fading channels. The noise coefficients $\boldsymbol{\rm{w}}\in\mathbb{C}^{K+1}$ are drawn from $\mathcal{CN}\left(0,\sigma_{n}^{2}\right)$, while the channel coefficient $h\in\mathbb{C}$ is drawn from $\mathcal{CN}\left(0,1\right)$ in the fading channel. To account for varying noise conditions, the noise variance $\sigma_{n}^{2}$ is computed for different $E_{\rm b}/N_{0}$. For simulations with a CFO, we draw $\psi$ from a uniform distribution $\pazocal{U}_{[0,2\pi)}$ and apply the transformation in (6) to $Y(z)$. In all simulations, we randomly sample $\boldsymbol{\rm{m}}$ from the set of all $2^{K}$ possible binary messages. Furthermore, we utilize the DiZeT decoder in all simulations and set $N=2^{10}$ for (13). The smooshing factor $\zeta$ is selected via a parameter sweep, based on minimizing BER in an AWGN channel with a CFO.
@@ IV-A Uncoded Error Rate Performance
We begin by comparing the performance of uncoded Huffman BMOCZ to uncoded SBMOCZ with $\zeta=0.0117$. For each simulation, we set $K=128$ and $\lambda=0.5$, and determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ. In Fig. 3 LABEL:sub@subfig:uncoded_performance_ber, we plot the BER for each scheme in AWGN and fading channels. Without a CFO, SBMOCZ performs roughly $1.3$ dB worse than Huffman BMOCZ in AWGN and $0.85$ dB worse in fading. However, when a CFO is present, Huffman BMOCZ fails (as indicated by the flat red curve), while SBMOCZ still functions but loses $1.46$ dB in AWGN and $2.92$ dB in fading relative to Huffman BMOCZ without a CFO. Observe that the BER curve with a CFO for SBMOCZ starts at a higher point compared to the version without a CFO, which occurs because the CFO estimate fails at lower $E_{\rm b}/N_{0}$, resulting in a cascade of bit errors. In Fig. 3 LABEL:sub@subfig:uncoded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We observe that SBMOCZ performs roughly the same across different CFO conditions, but is still approximately $1.5$ dB worse in AWGN and $1$ dB in fading as compared to Huffman BMOCZ without a CFO. Again, we observe that uncoded Huffman BMOCZ fails under a CFO.
@@ IV-B Coded Error Rate Performance
For these simulations, we utilize a (127,106)-BCH code for SBMOCZ. For Huffman BMOCZ, we employ a (127,106)-BCH absent a CFO and a (127,106)-ACPC under a CFO. To ensure fair comparison, the ACPC is implemented using a BCH code as the CPC. We note that (127,106)-BCH corrects up to three bit errors, while (127,106)-ACPC corrects a maximum of two, sacrificing one bit of error correction to detect the cyclic shift $m\Delta{\phi}$. For (127,106)-ACPC, we implement an inverse discrete Fourier transform (IDFT)-based DiZeT decoder with an oversampling factor of $Q=200$ to estimate and correct the fractional component $\psi_{0}$. Further details on the implementation of the ACPC can be found in [6]. In each scheme, we set $K=127$ and $\lambda=0.5$, using $\zeta=0.0130$ for SBMOCZ. Furthermore, we determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ.
In Fig. 4 LABEL:sub@subfig:coded_performance_ber, we plot the BER for each scheme in AWGN and fading channels. In the AWGN channel, SBMOCZ and ACPC perform similarly at higher $E_{\rm b}/N_{0}$, while ACPC performs much worse at lower $E_{\rm b}/N_{0}$. For a higher $E_{\rm b}/N_{0}$, both SBMOCZ and ACPC perform roughly $1.6$ dB worse relative to coded Huffman BMOCZ without a CFO. However, in the fading channel, SBMOCZ achieves a significant $4$ dB gain over ACPC. In Fig. 4 LABEL:sub@subfig:coded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We find that SBMOCZ shows a loss of around $0.65$ dB in AWGN and $0.6$ dB in fading relative to ACPC. Nevertheless, SBMOCZ demonstrates a large BER gain over ACPC in the fading channel.
@@ V Concluding Remarks
In this study, we introduce a modified BMOCZ zero constellation termed SBMOCZ in which the angular separation between adjacent zeros, excluding the first and last, is reduced. By smooshing the zeros closer together, we create a constellation gap that rotates under a CFO, which enables the receiver to estimate and correct the rotation by identifying the gap’s position. Compared to uncoded Huffman BMOCZ, we find that uncoded SBMOCZ works under a CFO, at the cost of a modest performance reduction absent a CFO. Against Huffman BMOCZ with ACPC, we find SBMOCZ shows a $4$ dB BER gain in a fading channel, with comparable performance in other scenarios. Future work will focus on optimization of SBMOCZ constellation parameters, such as the smooshing factor and the radii for each zero. Furthermore, we will explore alternative decoding approaches for SBMOCZ.
Appendix A Radius Derivation
@@ Appendix B Proof of Lemma 1
Let $\boldsymbol{\alpha}=(r_{\text{sb}}e^{j\phi_{0}},r_{\text{sb}}e^{j\phi_{1}},\dots,r_{\text{sb}}e^{j\phi_{K-1}})\in\mathbb{C}^{K}$ be the zeros of an SBMOCZ polynomial $X(z)$. Evaluating on the unit circle and taking the squared magnitude yields
$|X(e^{-j\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{K-1}\left|e^{-j\theta}-r_{\text{sb}}e^{j\phi_{k}}\right|^{2}$ (1b)
@@
$f(\theta,\phi_{k})=\left|e^{-j\theta}-r_{\text{sb}}e^{j\phi_{k}}\right|^{2}=1+r_{\text{sb}}^{2}-2r_{\text{sb}}\cos(\theta+\phi_{k})\;.$ (2b)
@@
$|X(e^{-j\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{\frac{K}{2}-1}f(\theta,\phi_{k})\prod_{k=0}^{\frac{K}{2}-1}f(\theta,-\phi_{k})\;.$ (3b)
@@
$\left|X(e^{-j\theta})\right|^{2}=\left|X(e^{j\theta})\right|^{2}\;.$ (4b)
@@
$|X(e^{-j\theta})|^{2}=|x_{K}|^{2}f(\theta,\pi)\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,\phi_{k})\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,-\phi_{k})\;.$ (5b)

### v3

@@ III-B CFO Correction with SBMOCZ
In this subsection, we introduce a method to correct a CFO impairment by evaluating the rotated polynomial $\tilde{Y}(z)$ at various points along the unit circle. To justify evaluating on the unit circle, we highlight some useful properties. We begin with the following proposition:
@@
Let $X(z)$ be a polynomial with the coefficient vector $\boldsymbol{\rm{x}}=[x_{0},x_{1},\dots,x_{K}]\in\mathbb{C}^{K+1}$. Then, the squared magnitude of $X({\rm e}^{-{\rm j}\theta})$ for $\theta\in\left[0,2\pi\right)$ corresponds to the discrete-time Fourier transform (DTFT) of the auto-correlation sequence $\mathbb{a}=[a_{-K},a_{-K+1},\dots,a_{K}]\in\mathbb{C}^{2K+1}$ for $\boldsymbol{\rm{x}}$. In particular, we have
$\left|X({\rm e}^{-{\rm j}\theta})\right|^{2}=\sum_{\ell=-K}^{K}a_{\ell}{\rm e}^{-{\rm j}\theta\ell}\;.$ (10)
@@
The auto-correlation sequence $\mathbb{a}\in\mathbb{C}^{2K+1}$ is the same for any transmitted BMOCZ coefficient sequence $\boldsymbol{\rm{x}}\in\mathbb{C}^{K+1}$ [4]. Consequently, for any SBMOCZ polynomial $X(z)$, the evaluation $|X({\rm e}^{-{\rm j}\theta})|^{2}$ yields the same result.
The important takeaway from Proposition 1 and Corollary 1 is that the smooshed zero constellation gap will induce a peak on the unit circle at the same location for any transmitted SBMOCZ coefficient sequence $\boldsymbol{\rm{x}}$. Additionally, evaluation on the unit circle will be symmetric about the peak, as highlighted by the following lemma:
@@
Let $X(z)$ be an SBMOCZ polynomial. When evaluating $|X({\rm e}^{-{\rm j}\theta})|^{2}$ with $\theta\in[0,2\pi)$, we have
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|X({\rm e}^{{\rm j}\theta})|^{2}\;.$ (11)
The proof is given in Appendix B. For SBMOCZ, the gap is centered on the positive real axis, resulting in a peak on the unit circle at $z=1$. To illustrate this behavior, we now present Fig. 2, which plots $|X(z)|^{2}$ along the unit circle for different $\zeta$ with $K=32$. As the gap increases in size, the peak becomes more pronounced, which makes the CFO estimate more robust against noise. However, this comes at the price of larger zero perturbation under noise. Additionally, Fig. 2 shows that Huffman BMOCZ is incompatible with our approach, since there is not a unique maximum on $\left[0,2\pi\right)$. Instead, evaluation on the unit circle exhibits sinusoidal behavior for Huffman BMOCZ. The key observation is that by identifying the peak location for $|\tilde{Y}(z)|^{2}$ on the unit circle, we can obtain an estimate for the angular rotation $\psi$.
Without loss of generality, we will now consider $|\tilde{Y}(z)|$, which peaks at the same location as its square. Under rotation, the maximum of $|X({\rm e}^{{\rm j}(\theta+\psi)})|$ in SBMOCZ occurs whenever $\theta=-\psi$, so we can estimate the angular rotation from $\tilde{Y}(z)$ via
$\hat{\psi}=\arg\!\!\!\,\max_{\theta\in\left[0,2\pi\right)}\left|\tilde{Y}({\rm e}^{-{\rm j}\theta})\right|\;.$ (12)
@@
$\hat{\psi}\approx\frac{2\pi}{N}\arg\max_{n\in[N]}\left|\tilde{Y}({\rm e}^{-{\rm j}2\pi n/N})\right|\;,$ (13)
where $N$ determines the resolution of the search space. We highlight that $\tilde{Y}({\rm e}^{-{\rm j}2\pi n/N})$ can be evaluated using a single $N$-point DFT of the coefficients, which results in a modest time complexity of $\mathcal{O}(N\log N)$ for the proposed algorithm. To compensate for the rotation, we perform the correction $\hat{Y}(z)=\tilde{Y}(z{\rm e}^{-{\rm j}\hat{\psi}})$, where $\hat{Y}(z)$ is an estimate of the received polynomial without a CFO impairment. The DiZeT decoder is then applied to $\hat{Y}(z)$ to recover the binary message $\boldsymbol{\rm{m}}\in\mathbb{Z}^{K}_{2}$.
@@ IV Numerical results
In this section, we compare the performance of SBMOCZ to Huffman BMOCZ in AWGN and flat-fading channels. The noise coefficients $\boldsymbol{\rm{w}}\in\mathbb{C}^{K+1}$ are drawn from $\mathcal{CN}\left(0,\sigma_{n}^{2}\right)$, while the channel coefficient $h\in\mathbb{C}$ is drawn from $\mathcal{CN}\left(0,1\right)$ for the fading channel. To account for varying noise conditions, the noise variance $\sigma_{n}^{2}$ is computed for different $E_{\rm b}/N_{0}$. For simulations using a CFO, we draw $\psi$ from the uniform distribution $\pazocal{U}_{[0,2\pi)}$ and apply the transformation in (6) to $Y(z)$. In each simulation, we randomly sample $\boldsymbol{\rm{m}}\in\mathbb{Z}^{K}_{2}$ from the set of all $2^{K}$ possible binary messages. Furthermore, we utilize the DiZeT decoder and set $N=1,024$ for the estimator in (13). The smooshing factor $\zeta$ is selected using a parameter sweep, based on minimizing BER under a CFO in AWGN.
@@ IV-A Uncoded Error Rate Performance
To begin, we compare the performance of uncoded Huffman BMOCZ to uncoded SBMOCZ using $\zeta=0.0117$. For each simulation, we set $K=128$ and $\lambda=1/2$, and determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ. In Fig. 3LABEL:sub@subfig:uncoded_performance_ber, we plot the BER curves for each scheme in AWGN and fading channels. Without a CFO, SBMOCZ performs roughly $1.3$ dB worse than Huffman BMOCZ in AWGN and $0.85$ dB worse in fading. However, when a CFO is introduced, Huffman BMOCZ fails, as indicated by the red curve, while SBMOCZ remains functional but loses $1.46$ dB in AWGN and $2.92$ dB in fading relative to Huffman BMOCZ without a CFO. Observe that the BER curve for SBMOCZ with a CFO starts at a higher point compared to the version without a CFO, which occurs because the CFO estimate fails at lower $E_{\rm b}/N_{0}$, resulting in a cascade of bit errors. In Fig. 3LABEL:sub@subfig:uncoded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We observe that SBMOCZ performs roughly the same with and without a CFO, but is still approximately $1.5$ dB worse in AWGN and $1$ dB in fading as compared to Huffman BMOCZ without a CFO. Again, we see that Huffman BMOCZ without channel coding does not function under a CFO.
@@ IV-B Coded Error Rate Performance
For these simulations, we utilize a (127,106)-BCH code for SBMOCZ. For Huffman BMOCZ, we employ a (127,106)-BCH code absent a CFO and a (127,106)-ACPC under a CFO. To ensure fair comparison, the ACPC is implemented using a BCH code as the CPC. We note that (127,106)-BCH corrects up to three bit errors, while (127,106)-ACPC only corrects a maximum of two, sacrificing one bit of error correction to detect the cyclic shift $m\Delta{\phi}$. For (127,106)-ACPC, we implement an inverse discrete Fourier transform (IDFT)-based DiZeT decoder with an oversampling factor of $Q=200$ to estimate and correct the fractional component $\psi_{0}$. Further details on the implementation of the ACPC can be found in [6]. In each scheme, we set $K=127$ and $\lambda=1/2$, using $\zeta=0.0130$ for SBMOCZ. Furthermore, we determine the radii using (5) for Huffman BMOCZ and (9) for SBMOCZ.
In Fig. 4LABEL:sub@subfig:coded_performance_ber, we plot the BER for each scheme in AWGN and fading channels. In AWGN, SBMOCZ and the ACPC perform similarly at higher $E_{\rm b}/N_{0}$, while the ACPC performs much worse at lower $E_{\rm b}/N_{0}$. For higher $E_{\rm b}/N_{0}$, both SBMOCZ and the ACPC perform roughly $1.6$ dB worse relative to coded Huffman BMOCZ without a CFO. However, in the fading channel, SBMOCZ achieves a significant $4$ dB gain over the ACPC. In Fig. 4LABEL:sub@subfig:coded_performance_bler, we plot the BLER for each scheme in AWGN and fading channels. We find that SBMOCZ shows a loss of around $0.65$ dB in AWGN and $0.6$ dB in fading compared to the ACPC. Nevertheless, SBMOCZ demonstrates a large BER gain over the ACPC in a fading channel.
@@ V Concluding Remarks
In this study, we introduce a new smooshed BMOCZ zero constellation called SBMOCZ in which the angular separation between adjacent zeros, excluding the first and last, is reduced. By smooshing the zeros closer together, we create a constellation gap that rotates under a CFO, which enables the receiver to estimate and correct the rotation by identifying the gap’s position. Compared to uncoded Huffman BMOCZ, we find that uncoded SBMOCZ functions under a CFO, at the cost of a modest performance reduction without a CFO. Against Huffman BMOCZ with an ACPC, we find that coded SBMOCZ achieves a $4$ dB BER gain in the fading channel, with comparable performance in other scenarios. Future work will focus on optimization of SBMOCZ constellation parameters, such as the smooshing factor and the radii for each zero. Moreover, we will explore alternative decoding approaches for SBMOCZ.
Appendix A Radius Derivation for SBMOCZ
@@ Appendix B Proof of Lemma 1
Let $\boldsymbol{\alpha}=[r_{\text{sb}}{\rm e}^{{\rm j}\phi_{0}},r_{\text{sb}}{\rm e}^{{\rm j}\phi_{1}},\dots,r_{\text{sb}}{\rm e}^{{\rm j}\phi_{K-1}}]\in\mathbb{C}^{K}$ be the zeros of an SBMOCZ polynomial $X(z)$. Evaluating on the unit circle and taking the squared magnitude yields
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{K-1}\left|{\rm e}^{-{\rm j}\theta}-r_{\text{sb}}{\rm e}^{{\rm j}\phi_{k}}\right|^{2}$ (1b)
@@
$f(\theta,\phi_{k})=\left|{\rm e}^{-{\rm j}\theta}-r_{\text{sb}}{\rm e}^{{\rm j}\phi_{k}}\right|^{2}=1+r_{\text{sb}}^{2}-2r_{\text{sb}}\cos(\theta+\phi_{k})\;.$ (2b)
@@
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|x_{K}|^{2}\prod_{k=0}^{\frac{K}{2}-1}f(\theta,\phi_{k})\prod_{k=0}^{\frac{K}{2}-1}f(\theta,-\phi_{k})\;.$ (3b)
@@
$\left|X({\rm e}^{-{\rm j}\theta})\right|^{2}=\left|X({\rm e}^{{\rm j}\theta})\right|^{2}\;.$ (4b)
@@
$|X({\rm e}^{-{\rm j}\theta})|^{2}=|x_{K}|^{2}f(\theta,\pi)\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,\phi_{k})\prod_{k=0}^{\lfloor\frac{K}{2}\rfloor-1}f(\theta,-\phi_{k})\;.$ (5b)